Light is travelling through air with an index of refraction of 1.00, at an angle of 30o, it then enters a mystery liquid and refracts at an angle of 14o This means that the index of refraction of the mystery liquid is:
1
Expert's answer
2020-12-21T11:00:36-0500

We can find the index of refraction of the mystery liquid from the Snell's Law:


"n_isin\theta_i=n_rsin\theta_r,"

here, "n_i=1.0" is the index of refraction of the incident medium (or the air), "n_r" is the index of refraction of the refractive medium (or the mystery liquid), "\theta_i=30^{\circ}" is the angle of incidence, "\theta_r=14^{\circ}" is the angle of refraction.

Then, from this equation we can calculate the index of refraction of the mystery liquid:


"n_r=\dfrac{n_isin\theta_i}{sin\theta_r}=\dfrac{1.0\cdot sin30^{\circ}}{sin14^{\circ}}=2.1"

Answer:

"n_r=2.1"